OpenStudy (anonymous):

Find the rule (quadratic equation) from this table (list of ordered pairs). (x, y), (-3, 6), (-2, 0), (-1, -4), (0, -6), (1, -6), (2, -4), (3, 0)

OpenStudy (anonymous):

I also have another question that is similar. So I want the method more than the answer. Any help is appreciated. : )

hero (hero):

First plot the points. I recommend using http://www.desmos.com to do so. From the points, try to figure out what the vertex (h,k) is. Then plug any two points into the following equations: y = a(x - h) + k y = a(x - h) + k and create a system of equations Then solve for a